OpenStudy (anonymous):

When we say \[\frac{\mathrm{d} y}{\mathrm{d} x}\], it also be written as \[\Delta x\] or \[\Delta y\]?

OpenStudy (anonymous):

not the same thing

OpenStudy (anonymous):

no thats different dy/dx is a derivatice delta y and delta x are changes

OpenStudy (anonymous):

Isn't the derivative the change? Like for dy/dx, it is the change in the y-axis for each x value? When what are delta x and delta y?

OpenStudy (anonymous):

the derivative can be written (in fact historically was almost always written) as \[\lim_{\Delta x \rightarrow 0}\frac{\Delta y}{\Delta x}\]

OpenStudy (amistre64):

put your hands 14 inches apart; \(\Delta\)x = 14 inches; as you bring your hands together that \(\Delta\)x gets smaller and smaller. The moment your hands touch and the distance is gone; you have reached dx

OpenStudy (amistre64):

the ghost of a departed value

OpenStudy (anonymous):

but notice it is a limit that gives you \[\frac{dy}{dx}\]as you take the limit, the greek letter becomes an english letter

OpenStudy (amistre64):

god save the queen!!

OpenStudy (anonymous):

sort of like \[\sum\] and \[\int\]

OpenStudy (anonymous):

who you calling a queen?

OpenStudy (amistre64):

♫♫...we are the champions .... ♫♫

OpenStudy (anonymous):

oh of course.

OpenStudy (amistre64):

what is the area of a circumference?

OpenStudy (anonymous):

So the delta x is the change in the x value, while the dy/dx is the rate of change, the gradient. Since the gradient is the rate of change, I remember the gradient is a function that tells the increase in the y-axis for every increment of x value on the x-axis, is this right? Then this gradient tells a change in the y-axis for each increment of x value, is delta y or delta x?